# Meta4XNLI: A Crosslingual Parallel Corpus for Metaphor Detection and Interpretation

### v1

@@ 2 Related Work
Metaphor interpretation datasets are more scarce and the lack of variety of languages is more striking than in metaphor detection corpora. The only multilingual dataset is MABL, however, data and annotations are not parallel between languages. We can find different datasets according to the tasks used to approach interpretation. Nevertheless, these corpora tend to bring along biases and artifacts as a result of synthetic annotation processes Boisson, Espinosa-Anke, and Camacho-Collados (2023). In datasets developed for paraphrase, tuples are commonly composed of a sentence with a metaphorical expression that is replaced for its literal meaning. In the case of NLI datasets, premise-hypothesis pairs are typically constructed by lexical substitution: the entailment is based on the literal paraphrase of the metaphor and contradiction on the replacement of the metaphorical terms by their antonym. Thus, we reckon these instances are not representative of spontaneous occurrences of metaphor in natural language utterances.
These shortcomings when developing datasets for figurative language can induce misleading results and conclusions. To bridge this gap, we bring forward the first parallel corpus that contains metaphorical annotations for metaphor detection at token level that cover all POS, and for metaphor interpretation via NLI in ES and EN. In addition, this resource includes a sufficient number of instances and contains data from natural language utterances that were not specifically generated for metaphor processing tasks. We summarised the distinct features of most popular available datasets in Table 1. We believe our proposed dataset is a valuable resource to continue with this line of research of multi- and cross-lingual approaches, which can allow to explore whether metaphors are transferable among languages.
Lang |
Task | Level | POS | Parallel | #sents | NL | |
@@
| IMPLI | EN | NLI | sent | - | ✗ | 1045 | ✗ |
| NewsMet | EN | paraphr | token | verb | ✗ | 1519 | ✓ |
| MABL |
HI, ID, JV, KN, SU, SW, YO |
NLI | sent | - | ✗ | 3709 | ✓ |
| Meta4XNLI |
ES, EN |
det, NLI | tok, sent | all | ✓ | 13320 | ✓ |
@@ 3 Meta4XNLI Corpus
To address these shortcomings, we introduce Meta4XNLI, a parallel dataset in ES and EN with metaphorical annotations for detection and interpretation via NLI in texts of natural language utterances and multiple domains. In the following subsections, we describe the collection of the dataset (Subsection 3.1), the methodology we employed to annotate metaphors for each task and language (Subsection 3.2) and, as a result, we will provide details of the outcoming dataset of this process (Subsection 3.3).
@@ 3.1 Description
Meta4XNLI is a compilation of XNLI Conneau et al. (2018) and esXNLI Artetxe, Labaka, and Agirre (2020). We decided to exploit these resources since we evaluate metaphor interpretation through NLI and these datasets were originally developed for this task. This enables both annotations at token level for detection and pair level for interpretation. In addition, it contains parallel text, from which we select data in ES and EN. Moreover, the combination of XNLI and esXNLI constitutes a dataset of large size compared to common available resources for metaphor processing and with natural language utterances and spontaneous usage of metaphors. The distribution of Meta4XNLI is detailed in Table 2.
XNLI dataset is a cross-lingual evaluation set for MultiNLI Williams, Nangia, and Bowman (2018). It contains parallel data with original text in EN subsequently human-translated to other 14 languages, among which we exploited only EN and ES. It comprehends a total of 7500 premise-hypothesis pairs from 10 text genres, that is, 830 premises and 2490 hypotheses from XNLIdev, and 1670 premises and 5010 hypotheses from XNLItest. esXNLI comprises a total of 2490 pairs, as in XNLIdev, from 5 different genres. In this case, sentences were originally collected in ES and then human-translated to EN. The direction of translation (EN > ES in XNLI and ES > EN in esXNLI) is an interesting feature to explore how translation may affect metaphor cross-lingual transfer and whether it impacts models performance.
The collection methodology is shared between XNLI and esXNLI: a set of premise sentences were crawled from various sources in EN and ES respectively. Afterwards, crowd-source workers were asked to generate three hypotheses for each premise, one for each label. Both corpora are balanced in terms of inference tags and text domains.
@@ 3.2 Annotation Process
The methodology to label Meta4XNLI varies across tasks and languages. In the following subsections, we will delve into the annotation processes used for detection and interpretation in each language. First, we developed ES annotations, from which we transferred a subset of detection labels and interpretation annotations in EN. Manual labeling was developed by a linguist native speaker of Spanish with advanced knowledge of English.
@@ 3.2.1 Detection
Annotations for this task are developed at token level, since we approach metaphor detection as a sequence labeling task. We extract premises and hypotheses sentences and annotate them separately. Therefore, we do not take into account the premise as context to annotate its corresponding hypotheses and vice versa. Due to this split, the total number of labeled sentences amounts to 13320. With respect to the type of metaphors, we consider as candidates the tokens belonging to a semantically significant part of speech (POS): nouns, verbs, adjectives and adverbs.
We adopt the MIPVU guidelines Steen et al. (2010) throughout the whole annotation process, either manual revision or in automatic predictions, since models used were trained on data labeled accordingly to this procedure. It can be summarised in four main steps:
@@
Read the entire text–discourse to establish a general understanding of the meaning.
@@
Determine the lexical units in the text–discourse.
@@
The annotation process in this language comprehends two phases, as depicted in Figure 2. First, we automatically label Meta4XNLIES by leveraging mDeBERTa He, Gao, and Chen (2021) fine-tuned for metaphor detection in ES with CoMeta. We choose mDeBERTa since it is the multilingual model that achieved highest F1 score in the experimental setup of Sanchez-Bayona and Agerri (2022). This choice aims at reducing the heavy work load and time investment that manual annotation from scratch requires. Afterwards, we manually inspect and correct the predictions in the whole dataset. From the first phase of automatic labeling, 748 tokens were predicted as metaphor in the premises and 724 in the hypotheses. After a complete and manual revision of all sentences, we removed 74 tokens and added 481 undetected metaphors in the premises; whereas in the hypotheses, we deleted 118 false positives and labeled 533 false negatives.
The main sources of ambiguity in ES emerge from multi-word expressions (MWE) and polysemy. The main issue when labeling MWE is to decide whether to treat the MWE as a single lexical unit, a fixed expression where the meaning of each word is not transparent anymore, or if it should be regarded as a collocation. Collocations are expressions where the constituent words tend to co-occur with high frequency but are not fixed, since each of its elements can be replaced by others with similar meaning.
@@
For instance, in Example 2, the expression entrar en pánico could be initially considered as a MWE with a single lexical unit, therefore, the three tokens would be labeled as metaphorical. However, this expression specifically means “to panic”, in which the verb entrar (lit. “to enter”) holds metaphorical meaning, since “panic” is not a physical place you can get into. In this case, we do not think of the expression as a fixed MWE, since the verb is also used metaphorically with other terms that are not places, like entrar en cólera (“to get angry”, lit. “to enter into wrath”) or entrar en calor (“to feel hot”, lit. “to enter into heat”). In all of these expressions the verb is conveying the sense of starting to feel the noun it complements, as if by entering to a place it transformed our sensitivity. The association of concepts arises from understanding emotions or sensations as physical locations.
Regarding polysemy, the existence of multiple and very nuanced senses associated to the same token can lead to confusion. It can be challenging to determine whether the basic meaning of the lexical unit is currently and generally known and used by native speakers or if they directly associate the lexical unit to the figurative meaning, not identifying the basic meaning at all.
@@
For instance, in Example 2 we label apoyo as metaphorical, since the most basic meaning of the verb apoyar in the dictionary defines it as “to make something rest upon another thing”. In this sentence, the contextual meaning refers to someone in favour of someone else’s goal. Figuratively, the goal can be understood in terms of a physical object so heavy that requires more than one anchor point to distribute its weight. Doubts in this examples come from the fact that the figurative sense can be more used that its basic one, thus speakers might not identify the metaphorical meaning as such. We make use of the Diccionario de la Real Academia Española (DRAE) as a tool to help us clarify these ambiguous cases. Nonetheless, metaphor identification remains a highly subjective task.
@@
We develop Meta4XNLIEN annotations semi-automatically and based on ES annotations, since our purpose is to publish a parallel resource to counterbalance monolingual English labeled data. These annotations serve as a starting point for further refinement and enable potential analysis of metaphors shared across these two languages. The whole process comprises mainly four phases, as depicted in Figure 3.
@@
The first step involves the projection of ES labels onto EN sentences. For this end, we utilised Easy-Label-Projection García-Ferrero, Agerri, and Rigau (2022), a tool developed for cross-lingual sequence labeling that makes use of word alignments and data- and model-transfer to project the labels from a source language (ES) to an untagged target language (EN). This mechanism is suitable when the labeled entity is certainly appearing in both source and target sentences, however, metaphors present in a source sentence are not necessarily in its translation. It depends on a series of multiple factors, namely the type of translation, the knowledge of the translator, if it is human-translated, or socio-cultural knowledge, among others. The next step comprises the manual revision of the projections.
In order to alleviate this issue, we made use of XLM-RoBERTa Conneau et al. (2020) trained on the VUAM dataset, as it is the multilingual model that showed best performance in the experiments in EN from the work of Sanchez-Bayona and Agerri (2022). Since projected sentences were manually reviewed (13% of the total), we only predicted automatically those sentences that did not receive a projection in the first place (87%). We manually reviewed the outcoming predictions from the MLM to correct errors and undetected metaphorical expressions following the MIPVU procedure. The number of metaphorical tokens extracted in each annotation phase is detailed in Table 3.
@@
There are some issues that emerged during the whole annotation process. Firstly, the projections phase entails that for a metaphorical expression to be annotated in EN, it has to have been annotated in ES in advance. Hence, metaphors in EN sentences that were not expressed figuratively in their ES counterpart will not be spotted, as in Example 3. In the ES sentence there is not a metaphorical expression annotated as such, however, in the EN version, the adjective heavy is holding metaphorical meaning, as a synonym of “demanding”, which is expressed literally in ES with the adjective exigentes. In this kind of cases where the metaphor is “gained in translation”, the lack of annotation in the source language implies no label will be projected onto the target sentence, missing a metaphorical instance. These examples give an account of how translation and the singularities of metaphors according to languages may affect to the annotation process of this task.
| Projections | Automatic | Total | ||||||
| ES | Lost | Gained | EN | Pred | Rev | EN | ||
| Premises | XNLIdev | 276 | 23 | 13 | 266 | 636 | 118 | 384 |
| XNLItest | 476 | 34 | 21 | 463 | 1364 | 216 | 679 | |
| esXNLI | 403 | 48 | 6 | 361 | 693 | 189 | 550 | |
| Hypotheses | XNLIdev | 276 | 10 | 33 | 299 | 1195 | 143 | 442 |
| XNLItest | 551 | 8 | 39 | 582 | 2437 | 282 | 864 | |
| esXNLI | 312 | 36 | 20 | 296 | 1199 | 115 | 411 | |
A los usuarios más exigentes se les debería cobrar más. The heavy users should be charged the most.
@@
Another issue that should be noted are false positives: all ES sentences with one or more labeled metaphors will transfer those tags to the EN sentence, regardless the translated tokens hold metaphoric meaning or not. To solve this question, we manually reviewed all sentences that received a projected tag. In this revision we had to remove metaphors that were “lost in translation” and adjust some spans projected to the target language, e.g. some annotated verbs in ES were projected in EN to the subject pronoun and verb, since some ES verbs forms are synthetic and include the person information in a single morpheme. Therefore, we eliminated the label from the pronoun and maintained only that of the verb. In Example 3 we can see how the verb peleaban (lit. “they fought”) labeled as metaphorical in the ES sentence (Example 3) was projected to the subject and verb in EN (Example 3). Example 3 represents the definitive version of the annotations after manual revision.
@@
Regarding the subset of sentences that were automatically labeled by XLM-RoBERTa, some concerns emerged with respect to annotations present in VUAM dataset during the phase of manual revision. Especially when it comes to phrasal verbs and abstract and polysemous terms that are lexicalised. For instance, due to Examples like 3 or 3 labeled as metaphorical in the training set VUAM, we observed a tendency to overannotate verbs that conform phrasal verbs as metaphorical, such as get, look, made or go. In most cases, these verbs appear in a context where they do not add any strong semantic information due to their lexicalization. Hence, we unmarked these instances that were predicted as metaphorical. Following this line of thought, we also discarded abstract and vague terms that are common-places of spontaneous discourse, like the word thing in Example 3, since it can refer to any kind of entity, either concrete or abstract and might not be directly matched to a more broadly used basic meaning.
@@ 3.2.2 Interpretation
Similar to other works cited in Section 2, we framed metaphor interpretation within the task of NLI Agerri (2008); Mohler, Tomlinson, and Bracewell (2013); Chakrabarty et al. (2021); Stowe, Utama, and Gurevych (2022); Kabra et al. (2023). Our approach aims at evaluating whether MLMs struggle to identify the relationship of inference when there is a metaphorical expression either in the premise, hypothesis or both sentences. To do so, we labeled premise-hypothesis pairs with metaphorical expressions where the understanding of the figurative expression is crucial to determine the inference label. We summarised this annotation process in the following steps:
@@
Establish the inference relation betweeen premise and hypothesis if not previously labeled.
@@
If there is any metaphorical expression either in premise or hypothesis:
@@
No: mark the pair as an non-relevant case.
@@
Premise: Hay tanto que se puede decir sobre eso, que sencillamente me voy a saltar eso. (lit. “There is so much you can say about that, that I am simply going to skip that”). Hypothesis: ¡Quiero contarte todo lo que sé sobre eso! (lit. “I want to tell you everything I know about it!”). Inference label: contradiction
Premise: No hay necesidad de hurgar en ese tema, a menos que quieras asegurarte de que nos hundimos. (lit. “There is no need to rummage in that topic, unless you want to make sure we will sink”). Hypothesis: Hay una forma de que se hundan. (lit. “There is one way to make them sink”). Inference label: entailment
Non-relevant cases comprehend premise-hypothesis pairs where the understanding of the literal sense of the metaphor is not essential to establish a relationship of entailment, contradiction or neutral. As we can see in Example 3.2.2, the metaphorical expression hurgar (lit. “to rummage”) is used with a sense of exploring an unpleasant topic, while the literal meaning implies to physically dig into an inner space. The entailment in this example is inferred from the likelihood of a sinking, not focusing on the willingness to talk about the mentioned topic. Thus, the interpretation of the metaphorical expression is not relevant to extract the inference relation.
We set aside the latter group as non-relevant cases, since we are not certain as to which extent the role of metaphor is significant in these occurrences. As a result, we discriminate three classes: a) pairs with metaphors that are relevant to the inference relationship, b) pairs with metaphors that are not relevant to the inference and c) pairs without metaphors.

### v2

@@ 2 Related Work
Datasets annotated for metaphor interpretation are even more scarce, and the lack of variety of languages is more pronounced than in metaphor detection corpora. The only multilingual dataset is MABL. However, neither the data nor the annotations are parallel across languages. Furthermore, while it is possible to find different datasets, these corpora tend to include biases or artifacts Boisson, Espinosa-Anke, and Camacho-Collados (2023). Thus, in paraphrasing datasets, tuples are commonly composed of a sentence with a metaphorical expression that is replaced with its literal meaning. In the case of NLI datasets, premise-hypothesis pairs are typically constructed by lexical substitution: the entailment is based on the literal paraphrase of the metaphor, whereas the contradiction is obtained by replacing the metaphorical expression with its antonym. Therefore, these instances are not representative of spontaneous occurrences of metaphor in natural language text.
| Dataset | Lang | Task | Level | POS | Parallel | #sents | NL |
@@
| IMPLI | EN | NLI | sent | - | ✗ | ∗1045 | ✗ |
| NewsMet | EN | paraphr | tok, sent | verb | ✗ | 1519 | ✓ |
| MABL | HI, ID, JV, KN, SU, SW, YO | NLI | sent | - | ✗ | 3709 | ✓ |
| Meta4XNLI | ES, EN | det, NLI | tok, sent | all | ✓ | 13320 | ✓ |
It can be argued that these shortcomings may produce misleading results and conclusions regarding the ability of language models to understand metaphorical language. In order to bridge this gap, we present the first parallel corpus including: (i) metaphorical annotations for metaphor detection at the token level covering nouns, adverbs, adjectives, and verbs, and (ii) metaphor interpretation annotations grounded in the NLI task for both ES and EN languages. In addition, this resource allows for large-scale cross-lingual and multilingual experimentation on both tasks (detection and interpretation) with data sourced from naturally occurring language, which was not synthetically generated via lexical substitution. Table 1 summarizes the main features of the most popular datasets for metaphor processing reviewed in this section.
@@ 3 Meta4XNLI Corpus
This section describes the development of Meta4XNLI, a parallel dataset in ES and EN newly annotated with metaphorical annotations for detection and interpretation. In the following subsections, we describe the collection of the dataset (Subsection 3.1), the methodology employed to annotate metaphor for each task and language (Subsection 3.2) and, finally, details of the resulting Meta4XNLI dataset (Subsection 3.3). We finish by discussing the inter-annotator agreement in Section 3.4 and properties of our dataset to study cross-lingual transfer of metaphorical expressions (Section 3.5).
@@ 3.1 Description
Meta4XNLI is a compilation of XNLI Conneau et al. (2018) and esXNLI Artetxe, Labaka, and Agirre (2020). We decided to use these data sources since we evaluate metaphor interpretation in the NLI framework. This enables both annotations at the token level for detection and sentence level for interpretation. In addition, it consists of parallel text, from which we select data in ES and EN. Moreover, the combination of XNLI and esXNLI constitutes a dataset of larger size compared to commonly available resources for metaphor processing, and with natural language utterances and spontaneous usage of metaphors. The distribution of Meta4XNLI is detailed in Table 2.
XNLI dataset is a cross-lingual evaluation set for MultiNLI Williams, Nangia, and Bowman (2018). It contains parallel data with original text in EN subsequently human-translated to other 14 languages, among which EN and ES were selected for manual annotation of metaphor. It includes a total of 7500 premise-hypothesis pairs from 10 text genres, namely, 830 premises and 2490 hypotheses from XNLIdev, and 1670 premises and 5010 hypotheses from XNLItest. esXNLI comprises a total of 2490 pairs, as in XNLIdev, from 5 different genres. In contrast to XNLI, sentences were originally collected in ES and then human-translated to EN. The direction of translation (EN > ES in XNLI and ES > EN in esXNLI) is an interesting feature to explore the cross-lingual transfer of metaphorical expressions.
XNLI and esXNLI share the same collection methodology: a set of premise sentences was crawled from various sources in EN and ES, respectively. Afterwards, crowd-source workers were asked to generate three hypotheses for each premise, one for each label. Both corpora are balanced in terms of inference tags and text domains.
@@ 3.2 Annotation Process
The methodology to label Meta4XNLI varies across tasks and languages. Although four annotators were involved in the whole process, manual annotation was mostly performed by a linguist native in Spanish with proficient knowledge of English and an expert in metaphor processing annotation.
@@ 3.2.1 Detection
Annotations for this task are performed at the token level, since we approach metaphor detection as a sequence labeling task. We extract premise and hypothesis sentences and annotate them separately. Therefore, we do not take into account the premise as context to annotate its corresponding hypotheses and vice versa. Due to this split, the total number of labeled sentences is 13320. With respect to the type of metaphors, we consider as candidates the tokens belonging to a semantically significant part of speech (POS): nouns, verbs, adjectives, and adverbs.
We adopt the MIPVU guidelines Steen et al. (2010) throughout the annotation process, either for manual revision or in automatic predictions, since models used were trained on data labeled accordingly to this procedure. The guidelines used can be summarised in the four main steps enumerated in Steen et al. (2010):
@@
Read the entire text/discourse to establish a general understanding of the meaning.
@@
Determine the lexical units in the text/discourse.
@@
The annotation process in this language comprehends two phases, as depicted in Figure 2. First, we automatically label Meta4XNLIES by leveraging mDeBERTa He, Gao, and Chen (2021) fine-tuned for metaphor detection in ES with CoMeta. We chose mDeBERTa since it is the multilingual model that achieved the highest F1 score in the experimental setup of Sanchez-Bayona and Agerri (2022). This choice reduces the heavy workload and time investment that a manual annotation from scratch would require. Afterwards, we manually inspect and correct the predictions for the whole dataset. From the first phase of automatic labeling, 748 tokens were predicted as metaphorical usages of words in the premises and 724 in the hypotheses. After a complete and manual revision of the 13320 sentences, 74 tokens were removed, and 481 undetected metaphors in the premises were added. Regarding the hypotheses, we deleted 118 false positives and labeled 533 false negatives.
The main sources of ambiguity in ES emerge from multi-word expressions (MWE) and polysemy. The main issue when labeling a MWE is to decide whether to treat the MWE as a single lexical unit, a fixed expression where the meaning of each word is not transparent anymore, or whether it should be regarded as a collocation. Collocations are expressions where the constituent words tend to co-occur with high frequency but are not fixed, since each of their elements can be replaced by others with similar meaning.
@@
For instance, in Example 2, the expression entrar en pánico could be initially considered as an MWE with a single lexical unit, therefore, the three tokens would be labeled as metaphorical. However, this expression specifically means “to panic”, in which the verb entrar (lit. “to enter”) holds metaphorical meaning, since “panic” is not a physical place you can get into. In this case, we do not think of the expression as a fixed MWE, since the verb is also used metaphorically with other terms that are not places, like entrar en cólera (“to get angry”, lit. “to enter into wrath”) or entrar en calor (“to feel hot”, lit. “to enter into heat”). In all of these expressions, the verb is conveying the sense of starting to feel the noun it complements, as if by entering a place, it transformed our sensitivity. This association of concepts arises from understanding emotions or sensations as physical locations.
Regarding polysemy, the existence of multiple and very nuanced senses associated with the same token can lead to confusion. It can be challenging to determine whether the basic meaning of the lexical unit is currently and generally known and used by native speakers or if they directly associate the lexical unit with the figurative meaning, not identifying the basic meaning at all.
@@
For instance, in Example 2 we label apoyo as metaphorical, since the most basic meaning of the verb apoyar in the dictionary defines it as “to make something rest upon another thing”. In this sentence, the contextual meaning refers to someone in favour of someone else’s goal. Figuratively, the goal can be understood in terms of a physical object so heavy that it requires more than one anchor point to distribute its weight. Doubts in this kind of examples arise from the fact that the figurative sense may be more frequent than its basic one, thus speakers might not identify the metaphorical meaning as such. We use the Diccionario de la Real Academia Española (DRAE) as a tool to help us clarify these ambiguous cases. Nonetheless, metaphor identification remains a subjective task.
@@
We generate Meta4XNLIEN annotations semi-automatically and based on ES annotations, since our purpose is to publish the parallel resource with annotations for metaphor detection and interpretation. The annotation process consists of four phases, as depicted in Figure 3.
@@
The first step involves the projection of ES labels onto EN sentences using Easy-Label-Projection García-Ferrero, Agerri, and Rigau (2022), a tool developed for cross-lingual sequence labeling that makes use of word alignments and data- and model-transfer to project the labels from a source language (ES) to an untagged target language (EN). This technique is appropriate when the labeled entity is certainly appearing in both source and target sentences. However, metaphors present in a source sentence are not necessarily in its translation. It depends on a series of multiple factors, namely the type of translation, the knowledge of the translator, whether it is human-translated, or socio-cultural knowledge, among others. The projected labels are manually revised.
The next step involves the automatic annotation (to minimize manual effort) using language models of the sentences without any metaphor annotation after executing the Spanish to English projection step (87% of the total). We applied XLM-RoBERTa Conneau et al. (2020) fine-tuned on the VUAM dataset (the best multilingual model in EN in the experiments by Sanchez-Bayona and Agerri (2022)). We manually reviewed the automatic predictions from XLM-RoBERTa to correct errors and undetected metaphorical expressions following the MIPVU procedure. The total number of metaphorical tokens is reported in the next Section 3.3.
@@
Some issues emerged during the whole annotation process. Firstly, the projections phase entails that for a metaphorical expression to be annotated in EN, it has to have been annotated in ES in advance. Hence, metaphors in EN sentences that were not expressed figuratively in their ES counterparts will not be spotted, as in Example 3. In the ES sentence, there is no metaphorical expression annotated as such. However, in the EN version, the adjective heavy is holding metaphorical meaning, as a synonym of “demanding”, which is expressed literally in ES with the adjective exigentes. In this type of case, the lack of annotation in the source language implies that no label will be projected onto the target sentence, missing a metaphorical instance. These examples give an account of how translation and the singularities of metaphors according to languages may affect the annotation process of this task. We provide an analysis of these cases in Section 3.5.
A los usuarios más exigentes se les debería cobrar más.
The heavy users should be charged the most.
@@
Another issue worth discussing is that of false positives: every ES sentence with one or more labeled metaphors will transfer those tags to the EN sentence, regardless of whether the translated tokens have metaphorical meaning or not. To address this, we manually reviewed all sentences with a projected metaphorical tag. This step removed metaphors that were “lost in translation” and adjusted some spans projected to the target language, e.g. some annotated verbs in ES were projected in EN to the subject pronoun and verb, since some ES verb forms are synthetic and include the person information in a single morpheme. Therefore, we eliminated the label from the pronoun and kept only that of the verb. In Example 3 we can see how the verb peleaban (lit. “they fought”) labeled as metaphorical in the ES sentence (Example 3) was projected onto the subject and verb in EN (Example 3). Example 3 represents the definitive version of the annotations after manual revision.
@@
Regarding the subset of sentences that were automatically labeled by XLM-RoBERTa, some concerns emerged with respect to annotations present in the VUAM dataset during the phase of manual revision. Especially when it comes to phrasal verbs and abstract and polysemous terms that are lexicalised. As illustrated by examples 3 and 3, which are labeled as metaphorical in the training set VUAM, we observed a tendency to overannotate verbs that compose phrasal verbs as metaphorical, such as get, look, made, or go. In most cases, these verbs appear in a context where they do not add any strong semantic information due to their lexicalisation. Hence, we unmarked these instances that were predicted as metaphorical. Following this line of thought, we also discarded abstract and vague terms that are common-places of spontaneous discourse, like the word thing in Example 3, since it can refer to any kind of entity, either concrete or abstract and might not be directly matched to a more broadly used basic meaning.
@@ 3.2.2 Interpretation
Similarly to other works discussed in Section 2, we frame metaphor interpretation within the task of NLI Agerri (2008); Mohler, Tomlinson, and Bracewell (2013); Chakrabarty et al. (2021); Stowe, Utama, and Gurevych (2022); Kabra et al. (2023). Our approach aims at evaluating the capability of LMs to identify the inferential relationship when there is a metaphorical expression either in the premise or, hypothesis, or in both sentences. To do so, we labeled premise-hypothesis pairs with metaphorical expressions where the understanding of the figurative expression is crucial to determine the inference label. We summarised this annotation process in the following steps:
@@
Establish the inference relation between premise and hypothesis if not previously labeled.
@@
If there is any metaphorical expression either in the premise or the hypothesis:
@@
No: mark the pair as a non-relevant case.
@@
Premise: Hay tanto que se puede decir sobre eso, que sencillamente me voy a saltar eso. (lit. “There is so much you can say about that, that I am simply going to skip that”).
Hypothesis: ¡Quiero contarte todo lo que sé sobre eso! (lit. “I want to tell you everything I know about it!”).
Inference label: contradiction
Premise: No hay necesidad de hurgar en ese tema, a menos que quieras asegurarte de que nos hundimos. (lit. “There is no need to rummage in that topic, unless you want to make sure we will sink”).
Hypothesis: Hay una forma de que se hundan. (lit. “There is one way to make them sink”).
Inference label: entailment
Non-relevant cases comprehend premise-hypothesis pairs where the understanding of the literal sense of the metaphor is not essential to establish a relationship of entailment, contradiction, or neutrality. As we can see in Example 3.2.2, the metaphorical expression hurgar (lit. “to rummage”) is used with a sense of exploring an unpleasant topic, while the literal meaning implies physically digging into an inner space. The entailment in this example is inferred from the likelihood of a sinking, not focusing on the willingness to talk about the mentioned topic. Thus, the interpretation of the metaphorical expression is not relevant to extracting the inference relation.
We set aside the latter group as non-relevant cases, since we are not certain as to which extent the role of metaphor is significant in these occurrences. As a result, we discriminate three classes: a) pairs with metaphors that are relevant to the inference relationship, b) pairs with metaphors that are not relevant to the inference, and c) pairs without metaphors.